# Topological characterization of magnon-polaron bands and thermal Hall conductivity in a frustrated kagome antiferromagnet

### v1

@@ II.1 Magnon Hamiltonian
$P_{1}=\frac{\sqrt{3}}{4}(1+\sin^{2}\eta),~{}~{}Q_{1}=\frac{1}{2}\sin{\eta},$
$R=-\frac{1}{4}+\frac{1}{4}\sin^{2}\eta-\frac{1}{2}\cos^{2}\eta,~{}~{}R_{1}=\frac{\sqrt{3}}{4}(1-\sin^{2}\eta),$
@@ II.2 Spin-phonon Hamiltonian
To incorporate the phonon environment, we now introduce the second term in Eq. (1), which corresponds to the phonon energy and is given by,
$H_{ph}=\hbar\omega\sum_{i}\tilde{b}_{i}^{\dagger}\tilde{b}_{i},$ (8)
where $\omega$ is the frequency of the longitudinal optical (LO) phonons and $\tilde{b}_{i}^{\dagger}(\tilde{b}_{i})$ denotes the phonon creation (annihilation) operator of the $i$-th lattice site. It is important to note that the LO frequency being reasonably high denotes the largest energy scale for our studies.
@@
$H_{sp}^{l}$ $=g^{l}\sum_{\langle i,j\rangle}\left(\tilde{b}_{i}^{\dagger}+\tilde{b}_{i}\right)\mathbf{S}_{i}\cdot\mathbf{S}_{j},$ (9a)
$H_{sp}^{nl}$ $=g^{nl}\sum_{\langle i,j\rangle}\left[\left(\tilde{b}_{i}^{\dagger}+\tilde{b}_{i}\right)-\left(\tilde{b}_{j}^{\dagger}+\tilde{b}_{j}\right)\right]\mathbf{S}_{i}\cdot\mathbf{S}_{j},$ (9b)
where $H_{sp}^{l}$ represents the local interaction with a single harmonic degree of freedom at a particular site $i$ that directly modifies the nearest-neighbour exchange integral with $g^{l}$ being the coupling strength. This type of spin-phonon coupling has been previously studied in the context of the spin-Peierls transition. In contrast, $H_{sp}^{nl}$ designates a non-local coupling, in which the nearest-neighbour exchange integral depends on the lattice spacing between the two neighbouring spins, or in other words, the difference between the phononic displacements ($\tilde{b}+\tilde{b}^{\dagger}$) separated by a NN lattice distance, where $g^{nl}$ is the corresponding coupling strength. As a result, this form of the spin-phonon coupling is more sensitive to the underlying lattice geometry than that of the local one ($g^{l}$). Due to this boson-boson interaction between the individual magnon excitation mode and the intrinsic phononic vibrational modes, the surrounding lattice of each magnon mode gets distorted (similar to the electron-phonon interaction in a fermionic lattice), giving rise to a polarization potential. Consequently, analogous to the polaronic scenario of an electronic system, a magnon-polaron quasiparticle bound state can be formed, which captures the essential physics of the resultant hybridized state without altering any underlying symmetry of the system.
@@ II.3 Magnon-phonon hybridization: canonical transformation approach
To form such a hybridized spectrum, we resort to the magnon generation by Eq. (5) for the $\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ terms in Eq. (9) and employ a canonical transformation thereafter with the generators $\small{R^{l}=\frac{g^{l}}{\hbar\omega}\sum_{\langle i,j\rangle}\left(\tilde{b}_{i}^{\dagger}-\tilde{b}_{i}\right)\mathbf{S}_{i}\cdot\mathbf{S}_{j}},$ and $\small{R^{nl}=\frac{g^{nl}}{\hbar\omega}\sum_{\langle i,j\rangle}\left[\left(\tilde{b}_{i}^{\dagger}-\tilde{b}_{i}\right)-\left(\tilde{b}_{j}^{\dagger}-\tilde{b}_{j}\right)\right]\mathbf{S}_{i}\cdot\mathbf{S}_{j}}$ for $H_{sp}^{l}$ and $H_{sp}^{nl}$, respectively, to decouple the magnon and phonon modes (for the detailed derivation see Appendix A.1). Subsequently, by taking a finite-temperature phonon averaging of the transformed Hamiltonian, $H$ in Eq. (1) can be reduced to $\tilde{H}_{R}$ with the renormalized factors containing coupling strengths (as derived in Appendix A.2). This decoupling scheme is extensively used in the context of electron-phonon interaction [72, 73] (in the form of a coherent Lang-Firsov generator), and also selectively in spin models via spin-Peierls transformation [95]. Therefore, we obtain a coherent state that is hybridized by magnon and phonons, where the corresponding renormalized Hamiltonian in the $k$-space, named magnon-polaron Hamiltonian, can be written in Bogoliubov–de Gennes form using the hybridized quasiparticle basis $(\Psi^{a}_{\mathbf{k}},\Psi^{b}_{\mathbf{k}},\Psi^{c}_{\mathbf{k}},\Psi^{a\dagger}_{-\mathbf{k}},\Psi^{b\dagger}_{-\mathbf{k}},\Psi^{c\dagger}_{-\mathbf{k}})^{T}$ as
$\tilde{H}_{R}(\mathbf{k})=\begin{bmatrix}A(\mathbf{k})&B(\mathbf{k})\\ B^{*}(-\mathbf{k})&A^{*}(-\mathbf{k})\end{bmatrix}$ (10)
@@
and all the components in $A(\mathbf{k})$ and $B(\mathbf{k})$ are given below as (see Appendix A),
@@
where $l$ and $nl$ symbolize cases where the system exhibits local and non-local magnon-phonon interactions, respectively. Moreover, all the symbols and parameters appearing in this expression are defined in Appendix A. Further, $\tilde{H}_{R}(\mathbf{k})$ follows generalized eigenvalue problem. Hence, to solve it, we introduce a non-Hermitian matrix $\tilde{H}^{\prime}_{R}(\mathbf{k})=\sigma_{3}\tilde{H}_{R}(\mathbf{k})$, where $\sigma_{3}=\sigma_{z}\otimes I_{3}$ is a metric, with $\sigma_{z}$ being the $z$-component of the Pauli matrices and $I_{3}$, an identity matrix. Similar to the magnons, the magnon-polaron hybridized states also obey the Bose-Einstein (BE) statistics. Further, we only focus on the three positive energy spectra among the possible six of them.
In the following discussion, we consider both types of couplings separately. Usually, the coherent decoupling method mentioned above works well in the high-frequency (anti-adiabatic) regime as the interaction deals with the LO phonons. However, at very large values of frequency, the terms $\Delta^{l(nl)}$ in Eq. (12) become dominant and lead to imaginary eigenvalues in the energy bands for both coupling mechanisms, indicating magnon-polaron instability in the corresponding ground state. Hence, to compare their respective effects on the magnon band topology in a consistent manner in our study, we fix the phonon energy scale as $\hbar\omega\gtrsim J_{1}S$, ensuring that $\hbar\omega\not\ll J_{1}S$ (thereby strictly avoiding the adiabatic regime). This restricts us to perform subsequent analyses in the aforementioned range of $\omega$.
@@ III Numerical analyses of magnon-polaron topology on KAFM
As mentioned earlier, due to the presence of extensive degeneracy in the energy bands, a frustrated kagome antiferromagnet is a strong candidate for a QSL phase. However, in the presence of $J_{2}>0$ and a nonzero out-of-plane DMI ($D\neq 0$), the degeneracy can be lifted, stabilizing the $\bm{Q}=0$, $120^{\circ}$ coplanar spin ordering LRO phase, as given in Eq. (3). Additionally, applying an out-of-plane magnetic field causes the spins to cant toward the field, resulting in the emergence of topological magnon bands and non-zero thermal Hall conductivity [20]. Further, in such a system, strong spin–phonon interaction involving optical phonons, whose energy is comparable to the exchange interaction strength $J_{1}$ can give rise to magnon-polaron hybridized quasiparticle states that retain the topological characteristics of the system. Moreover, we shall show that, without disturbing the classical ground state described by Eq. (3), the spin–phonon coupling strength $g^{l(nl)}$ can drive topological phase transitions associated with these hybridized states. Additionally, its interplay with the out-of-plane DMI produces interesting results.
As stated earlier, the presence of strong spin–phonon coupling may lead to magnon-polaron instability at a larger frequency $\omega>\omega_{c}$ ($\omega_{c}$ being the critical phonon frequency). To avoid this issue, even though the value of $\omega_{c}$ differs in the two cases corresponding to local and non-local spin-phonon couplings, we fix $\hbar\omega=J_{1}S$ throughout this study for consistency. Furthermore, we shall elaborate on the rationale for choosing $\omega$ using a band analogy in the following discussion.
III.1 Bulk spectra and topological transition by magnon-phonon coupling
@@
Figure 2: The bulk band structures are shown at the symmetry points inside the Brillouin zone. In the absence of any spin-phonon coupling, the bulk band structures for (a) $B_{0}=D=0$, $J_{2}=0.2J_{1}$ and (b) $B_{0}=0.4B_{s}$, $J_{2}=D=0.03J_{1}$ are plotted. (c) In the presence of local spin-phonon coupling with $\hbar\omega=J_{1}S$, $g^{l}=0.2J_{1}$, $B_{0}=0.4B_{s}$, $J_{2}=D=0.03J_{1}$ the bulk band structure is shown at zero temperature. In (b) and (c), the yellow-highlighted numbers correspond to the Chern numbers. (d) A combined phase diagram based on the Chern number sets of corresponding bands is shown as a function of $g^{l}$ and $D$.
@@
While our primary motivation is to explore the topological behavior of magnon-polaron states and its effect on the thermal Hall conductivity, for the sake of a comprehensive understanding, it is first important to discuss the origin of the LRO state in such a frustrated kagome antiferromagnet, as well as the key elements that enable the system to exhibit Dirac semimetal behavior. In the absence of any long-range interactions and DMI, the energy band structure obtained from Eq.(2) for the $120^{\circ}$ coplanar spin configuration exhibits zero-energy modes throughout the Brillouin zone (BZ), indicating a gapless spectrum between the ground and the excited states. However, upon introducing NNN interaction with $J_{2}>0$, the bands become dispersive and acquire finite energy along the high-symmetry points inside the BZ, such as, $\mathbf{\Gamma}=(0,0)$, $\mathbf{M}=(\frac{\pi}{2},\frac{\pi}{2\sqrt{3}})$ and $\mathbf{K}=(\frac{2\pi}{3},0)$ as seen in Fig. 2(a). Since the system preserves valley symmetry [98], the energies at the two Dirac points $\mathbf{K}$ and $\mathbf{K^{\prime}}$ remain degenerate. Also, in Fig. 2(a), all the three energy bands exhibit Goldstone modes at the $\mathbf{\Gamma}$ point due to the presence of SO(3) rotational symmetry, where it is reduced to SO(2) symmetry due to the presence of an out-of-plane DMI, resulting in only one Goldstone mode occuring at the $\mathbf{\Gamma}$ point. Furthermore, the middle and the upper bands cross linearly at the Dirac point $\mathbf{K}$, which makes the system a Dirac semimetal.
@@
Since the TRS operator $\tau$ flips the spin direction, any magnetically ordered system violates TRS and instead follows an effective TRS. Here, for this $120^{\circ}$ Néel order, the effective TRS is the combination of TRS and the mirror symmetry with respect to the kagome plane [20]. While out-of-plane DMI could not break this symmetry, an external magnetic field ($B_{0}$) causes the spins to be canted along the field direction, which breaks effective TRS. Further, from Eqs. (12) and (7), it can be shown that in the absence of an external magnetic field, as long as $\eta=0$, the $\mathbf{k}$-independent imaginary part of the hopping strength remains zero. However, for a nonzero value of $\eta$, the Hamiltonian breaks time-reversal symmetry, that is, $\widetilde{H}_{R}(-\mathbf{k})\neq\widetilde{H}_{R}^{*}(\mathbf{k})$. Consequently, in the presence of both $D$ and $B_{0}$, we observe gapped magnons everywhere inside the BZ with one Goldstone mode occurring at $\mathbf{\Gamma}$, as shown in Fig. 2(b).
@@
Further, to determine whether these band gaps are topological or trivial in such a TRS-broken system, one can compute the Chern numbers of the corresponding bands. The Chern number ($C_{n}$) associated with $n^{\text{th}}$ band is given by [99],
@@
where $n$ is the band index and $\Omega_{n}^{z}(\mathbf{k})$ is the Berry curvature corresponding to the $n^{\text{th}}$ band. Moreover, in the presence of broken TRS, Berry curvature becomes nonzero and is given by [26, 73],
@@
where $\psi_{n}(\mathbf{k})$ is the eigenfunction corresponding to the $n^{\text{th}}$ band. In general, the Berry curvature attains larger values near the high symmetry points ($\mathbf{\Gamma}$, $\mathbf{M}$, $\mathbf{K}$), where the dispersive bands exhibit a large curvature. Therefore, upon calculating the Chern number from Eq. (13), we find that the Chern numbers for the lower, middle, and upper bands are $3$, $-4$, and $1$, respectively, as shown in Fig. 2(b).
@@
Here, from Eq. (12), we observe that in the presence of local spin-phonon interaction ($H_{sp}^{l}$), the effective hopping strengths are reduced by a factor of $e^{-\lambda_{q}^{l}}$, where $q\in 1,2,3,4$ (see Eq. (32)), which helps to stabilize the low-energy magnon-polaron states. On the other hand, the onsite energy is modified by a negative energy shift $\Delta^{l}$ proportional to $(\alpha^{l})^{2}\omega$, which tends to destabilize the magnon-polaron states. Since our system already hosts a Goldstone mode near the $\mathbf{\Gamma}$ point, if the effect of $\Delta^{l}$ outweighs that of the $e^{-\lambda_{q}^{l}}$ term at large $\omega$, the system may slip into a regime with magnon-polaron instability, leading to imaginary eigenvalues within the BZ. Therefore, we keep $\hbar\omega=J_{1}S$, where the exponential factor dominates and hence avoids the magnon-polaron instability.
Figure 3: The bulk band structures are displayed along the high symmetry path, corresponding to the local spin-phonon coupling for $B=0.4B_{s}$, $J_{2}=0.03J_{1}$, $D=0.045J_{1}$ with (a) $g^{l}=0.125J_{1}$, (b) $g^{l}=0.3J_{1}$, (c) $g^{l}=0.586J_{1}$, (d) $g^{l}=0.8J_{1}$ at zero temperature. The yellow-highlighted numbers denote the Chern numbers corresponding to the bands shown in (b) and (c).
Figure 4: (a) A combined phase diagram based on the Chern number sets of the corresponding bands is depicted as a function of non-local coupling $g^{nl}$ and $D$. The bulk band structures are shown along the high symmetry points with $B=0.4B_{s}$, $J_{2}=0.03J_{1}$, $D=0.045J_{1}$ for (b) $g^{nl}=0.3J_{1}$, (c) $g^{nl}=0.438J_{1}$, (d) $g^{nl}=0.8J_{1}$ at zero temperature. The yellow-highlighted numbers denote the Chern numbers corresponding to the bands.
@@
In addition, to ensure that the system undergoes a topological phase transition, we calculate the Chern numbers corresponding to these bands and observe a transition from $C=3\rightarrow C=1$ for the lower band and from $C=-4\rightarrow C=-2$ for the middle band, as shown in Fig. 2(c). Hence, we observe that the presence of local spin-phonon coupling can significantly influence the topological state of the system. Furthermore, to examine the interplay between DMI and the spin-phonon coupling strength, we construct a topological phase diagram, shown in Fig. 2(d). Here, distinct topological regions are denoted by the Chern number set $C=(C_{1},C_{2},C_{3})$, where the indices $1$, $2$, and $3$ correspond to the lower, middle, and upper bands, respectively. Furthermore, in this phase diagram, we fix the other parameters at values given by $B=0.4J_{1}$, $J_{2}=0.03J_{1}$, and set the temperature to zero ($T=0$). Then, by assuring $D$ to be fixed and varying $g^{l}$ from $1$ to $J_{1}$, we observe multiple topological phase transitions along the $\text{u}_{1}$ and $\text{u}_{2}$ transition lines corresponding to a given classical ground state. Moreover, each of these transitions must be associated with a gap-closing transition, as can be seen from spectral features.

### v2

@@ II.1 Magnon Hamiltonian
$P_{1}=\frac{\sqrt{3}}{4}\left(1+\sin^{2}\eta\right),~{}~{}Q_{1}=\frac{1}{2}\sin{\eta},$
$R=-\frac{1}{4}+\frac{1}{4}\sin^{2}\eta-\frac{1}{2}\cos^{2}\eta,~{}~{}R_{1}=\frac{\sqrt{3}}{4}\left(1-\sin^{2}\eta\right),$
@@ II.2 Spin-phonon Hamiltonian
To incorporate the phonon environment, we now introduce the second term in Eq. (1), which corresponds to the phonon energy and is given by
$H_{\text{ph}}=\hbar\omega\sum_{i}\tilde{b}_{i}^{\dagger}\tilde{b}_{i},$ (8)
where $\omega$ is the frequency of the longitudinal optical (LO) phonons and $\tilde{b}_{i}^{\dagger}$ $(\tilde{b}_{i})$ denotes the phonon creation (annihilation) operator of the $i$-th lattice site. It is important to note that, as the LO frequency is reasonably high, it denotes the largest energy scale for our studies.
@@
$H_{\text{sp}}^{l}$ $=g^{l}\sum_{\langle i,j\rangle}\left(\tilde{b}_{i}^{\dagger}+\tilde{b}_{i}\right)\mathbf{S}_{i}\cdot\mathbf{S}_{j},$ (9a)
$H_{\text{sp}}^{nl}$ $=g^{nl}\sum_{\langle i,j\rangle}\left[\left(\tilde{b}_{i}^{\dagger}+\tilde{b}_{i}\right)-\left(\tilde{b}_{j}^{\dagger}+\tilde{b}_{j}\right)\right]\mathbf{S}_{i}\cdot\mathbf{S}_{j},$ (9b)
where $H_{\text{sp}}^{l}$ represents the local interaction with a single harmonic degree of freedom at a particular site $i$ that directly modifies the nearest-neighbour exchange integral with $g^{l}$ being the coupling strength. This type of spin-phonon coupling has been previously studied in the context of the spin-Peierls transition. In contrast, $H_{\text{sp}}^{nl}$ designates a non-local coupling, in which the NN exchange integral depends on the lattice spacing between the two neighbouring spins, or in other words, the difference between the phononic displacements ($\tilde{b}+\tilde{b}^{\dagger}$) separated by a NN lattice distance, where $g^{nl}$ is the corresponding coupling strength. As a result, this form of the spin-phonon coupling is more sensitive to the underlying lattice geometry than that of the local one ($g^{l}$). Due to this boson-boson interaction between the individual magnon excitonic mode and the intrinsic phononic vibrational mode, the surrounding lattice of each magnon mode gets distorted (similar to the electron-phonon interaction in a fermionic lattice), giving rise to a polarization potential. Consequently, analogous to the polaronic scenario of an electronic system, a magnon-polaron quasiparticle bound state can be formed, which captures the essential physics of the resultant hybridized state without altering any underlying symmetry of the system.
@@ II.3 Magnon-phonon hybridization: canonical transformation approach
To form such a hybridized spectrum, we resort to the magnon generation by Eq. (5) for the $\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ terms in Eq. (9) and employ a canonical transformation thereafter with the generators $\small{R^{l}=\frac{g^{l}}{\hbar\omega}\sum_{\langle i,j\rangle}\left(\tilde{b}_{i}^{\dagger}-\tilde{b}_{i}\right)\mathbf{S}_{i}\cdot\mathbf{S}_{j}},$ and $\small{R^{nl}=\frac{g^{nl}}{\hbar\omega}\sum_{\langle i,j\rangle}\left[\left(\tilde{b}_{i}^{\dagger}-\tilde{b}_{i}\right)-\left(\tilde{b}_{j}^{\dagger}-\tilde{b}_{j}\right)\right]\mathbf{S}_{i}\cdot\mathbf{S}_{j}}$ for $H_{\text{sp}}^{l}$ and $H_{\text{sp}}^{nl}$, respectively, to decouple the magnon and phonon modes (for the detailed derivation see Appendix A.1). Subsequently, by taking a finite-temperature phonon averaging of the transformed Hamiltonian, $H$ in Eq. (1) can be reduced to $\widetilde{H}_{R}$ with the renormalized factors containing coupling strengths (as derived in Appendix A.2). This decoupling scheme is extensively used in the context of electron-phonon interaction [72, 73] (in the form of a coherent Lang-Firsov generator), and also selectively in spin models via spin-Peierls transformation [95]. Therefore, we obtain a coherent state that is hybridized by magnon and phonons, where the corresponding renormalized Hamiltonian in the $k$-space, named magnon-polaron Hamiltonian, can be written in Bogoliubov–de Gennes form using the hybridized quasiparticle basis $(\Psi^{a}_{\mathbf{k}},\Psi^{b}_{\mathbf{k}},\Psi^{c}_{\mathbf{k}},\Psi^{a\dagger}_{-\mathbf{k}},\Psi^{b\dagger}_{-\mathbf{k}},\Psi^{c\dagger}_{-\mathbf{k}})^{T}$ as
$\widetilde{H}_{R}(\mathbf{k})=\begin{bmatrix}A(\mathbf{k})&B(\mathbf{k})\\ B^{*}(-\mathbf{k})&A^{*}(-\mathbf{k})\end{bmatrix},$ (10)
@@
and all the components in $A(\mathbf{k})$ and $B(\mathbf{k})$ are given below as (see Appendix A)
@@
where $l$ and $nl$ symbolize cases where the system exhibits local and non-local magnon-phonon interactions, respectively. Moreover, all the symbols and parameters appearing in this expression are defined in Appendix A. Further, $\widetilde{H}_{R}(\mathbf{k})$ follows generalized eigenvalue problem. Hence, to solve it, we introduce a non-Hermitian matrix $\widetilde{H}^{\prime}_{R}(\mathbf{k})=\sigma_{3}\widetilde{H}_{R}(\mathbf{k})$, where $\sigma_{3}=\sigma_{z}\otimes I_{3}$ is a metric, with $\sigma_{z}$ being the $z$-component of the Pauli matrices and $I_{3}$, an identity matrix. Similar to the magnons, the magnon-polaron hybridized states also obey the Bose-Einstein (BE) statistics. Further, we only focus on the three positive energy spectra among the possible six of them.
In the following discussion, we consider both types of couplings separately. Usually, the coherent decoupling method mentioned above works well in the high-frequency (anti-adiabatic) regime as the interaction potential deals with the LO phonons. However, at sufficiently large frequencies, the term $\Delta^{l(nl)}$ in Eq. (12) becomes dominant and leads to imaginary eigenvalues in the energy bands for both the coupling mechanisms, which eventually indicates that the initial ground state is no longer an energy minimum state. This is an artefact of the strong spin-phonon coupling, where the LO phonons significantly affect the renormalized energy bands, and hence have a considerable impact on the classical ground state. Thus, depending on the coupling strength ($g^{l(nl)}$) and phonon frequency ($\omega$), the ground state may either be stable or unstable, while the latter scenario is termed as magnon-polaron instability.
@@ III Numerical analyses of magnon-polaron topology on KAFM
As mentioned earlier, due to the presence of extensive degeneracy in the energy bands, a frustrated kagome antiferromagnet is a strong candidate as a QSL. However, in the presence of $J_{2}>0$ and a nonzero out-of-plane DMI ($D\neq 0$), the degeneracy can be lifted, stabilizing the $\bm{Q}=0$, $120^{\circ}$ coplanar spin ordering (as given in Eq. (3)), which manifests an LRO phase. Additionally, applying an out-of-plane magnetic field causes the spins to cant toward the field, resulting in the emergence of topological magnon bands and non-zero thermal Hall conductivity [20]. Further, in such a system, strong spin–phonon interaction involving optical phonons, whose energy is comparable to the exchange interaction strength $J_{1}$ can give rise to magnon-polaron hybridized quasiparticle states that retain the topological characteristics of the system. Moreover, we shall show that, without disturbing the classical ground state described by Eq. (3), the spin–phonon coupling strength $g^{l(nl)}$ can drive topological phase transitions associated with these hybridized states. Additionally, its interplay with the out-of-plane DMI produces interesting results.
As stated earlier, the presence of strong spin–phonon coupling may lead to magnon-polaron instability at larger frequencies, that is $\omega>\omega_{c}$ ($\omega_{c}$ being the critical phonon frequency). Hence, to avoid this issue, we need to analyze the magnitude of the critical frequency. Since, the band energy is being modified due to the effects of spin-phonon coupling strength and the phonon frequency (see Appendix A), we numerically estimate the corresponding values of critical phonon frequency $\omega_{c}$ (above which the instability occurs) as a function of the coupling strength $g^{l(nl)}$, while keeping all the other parameters fixed at $B_{0}=0.4B_{s}$, $J_{2}=0.03J_{1}$, $D=0.045J_{1}$, $T=0$, as shown in Fig. 2. These parameter values are employed in most of the subsequent analyses, and it should be noted that $\omega_{c}$ is insensitive to minor changes in the choices of $J_{2}$ and $D$. Additionally, we observe that the non-local coupling mechanism is more sensitive to the phonon frequencies, as considerably lower values of $\omega_{c}$ can induce instability, as compared to the local one. Hence, we fix phonon energy as $\hbar\omega=J_{1}S$ throughout this study for consistency, as for both the coupling mechanisms $\hbar\omega_{c}$ is greater than $J_{1}S$. Moreover, this particular value of $\hbar\omega$ (in fact any value in its vicinity) simultaneously aids in avoiding the adiabatic regime (maintaining $\hbar\omega\not\ll J_{1}S$) and eliminates the possibility of magnon-polaron instability.
Figure 2: The behavior of the critical frequency ($\omega_{c}$) as a function of both local and non-local coupling strengths $g^{l(nl)}$ is shown, with $\widetilde{\omega}_{c}=\hbar\omega_{c}/J_{1}S$. The other parameters are fixed at $B_{0}=0.4B_{s}$, $J_{2}=0.03J_{1}$, $D=0.045J_{1}$, $T=0$.
III.1 Bulk spectra and topological transition by spin-phonon coupling
@@
Figure 3: The bulk band structures are shown at the symmetry points inside the Brillouin zone. In the absence of any spin-phonon coupling, the bulk band structures for (a) $B_{0}=D=0$, $J_{2}=0.2J_{1}$ and (b) $B_{0}=0.4B_{s}$, $J_{2}=D=0.03J_{1}$ are plotted. (c) In the presence of local spin-phonon coupling with $\hbar\omega=J_{1}S$, $g^{l}=0.2J_{1}$, $B_{0}=0.4B_{s}$, $J_{2}=D=0.03J_{1}$ the bulk band structure is shown at zero temperature. In (b) and (c), the numbers highlighted in yellow correspond to the Chern numbers. (d) A combined phase diagram based on the Chern number sets of corresponding bands is depicted in the $g^{l}$-$D$ plane, where the topological transitions are demonstrated explicitly.
@@
While our primary motivation is to explore the topological behavior of magnon-polaron states and its effect on the thermal Hall conductivity, for the sake of a comprehensive understanding, it is first important to discuss the origin of the LRO phase in such a frustrated kagome antiferromagnet, as well as the key elements that enable the system to exhibit Dirac semimetal behavior. In the absence of any long-range interactions and DMI, the energy band structure obtained from Eq. (2) for the $120^{\circ}$ coplanar spin configuration exhibits zero-energy modes throughout the Brillouin zone (BZ), indicating a gapless spectrum between the ground and the excited states. However, upon introducing NNN interaction with $J_{2}>0$, the bands become dispersive and acquire finite energy along the high-symmetry points inside the BZ, such as, $\mathbf{\Gamma}=(0,0)$, $\mathbf{M}=(\frac{\pi}{2},\frac{\pi}{2\sqrt{3}})$ and $\mathbf{K}=(\frac{2\pi}{3},0)$ as seen in Fig. 3(a). Since the system preserves valley symmetry [98], the energies at the two Dirac points $\mathbf{K}$ and $\mathbf{K^{\prime}}$ remain degenerate. Also, in Fig. 3(a), all the three energy bands exhibit Goldstone modes at the $\mathbf{\Gamma}$ point due to the presence of SO(3) rotational symmetry, where it is reduced to SO(2) symmetry due to the presence of an out-of-plane DMI, resulting in only one Goldstone mode occuring at the $\mathbf{\Gamma}$ point [20]. Furthermore, the middle and the upper bands cross linearly at the Dirac point $\mathbf{K}$, which makes the system a Dirac semimetal.
@@
Since the TRS operator $\tau$ flips the spin direction, any magnetically ordered system violates TRS and instead follows an effective TRS. Here, for this $120^{\circ}$ Néel order, the effective TRS is the combination of TRS and the mirror symmetry with respect to the kagome plane. While out-of-plane DMI could not break this symmetry, an external magnetic field ($B_{0}$) causes the spins to be canted along the field direction, which breaks effective TRS. Further, from Eqs. (7) and (12), it can be shown that in the absence of an external magnetic field, as long as $\eta=0$, the $\mathbf{k}$-independent imaginary part of the hopping strength remains zero. However, for a nonzero value of $\eta$, the Hamiltonian breaks time-reversal symmetry, that is, $\widetilde{H}_{R}(-\mathbf{k})\neq\widetilde{H}_{R}^{*}(\mathbf{k})$. Consequently, in the presence of both $D$ and $B_{0}$, we observe gapped magnons everywhere inside the BZ with one Goldstone mode occurring at $\mathbf{\Gamma}$, as shown in Fig. 3(b).
@@
Further, to determine whether these band gaps are topological or trivial in such a TRS-broken system, one can compute the Chern numbers of the corresponding bands. The Chern number ($C_{n}$) associated with $n^{\text{th}}$ band is given by [99]
@@
where $n$ is the band index and $\Omega_{n}^{z}(\mathbf{k})$ is the Berry curvature corresponding to the $n^{\text{th}}$ band. Moreover, in the presence of broken TRS, Berry curvature becomes nonzero and is given by [26, 73]
@@
where $\psi_{n}(\mathbf{k})$ is the eigenfunction corresponding to the $n^{\text{th}}$ band. In general, the Berry curvature attains larger values near the high symmetry points ($\mathbf{\Gamma}$, $\mathbf{M}$, $\mathbf{K}$, $\mathbf{K^{\prime}}$), where the dispersive bands exhibit a large curvature. Therefore, upon calculating the Chern number from Eq. (13), we find that the Chern numbers for the lower, middle, and upper bands are $3$, $-4$, and $1$, respectively, as shown in Fig. 3(b).
Figure 4: The bulk band structures are displayed along the high symmetry path, corresponding to the local spin-phonon coupling for $B=0.4B_{s}$, $J_{2}=0.03J_{1}$, and $D=0.045J_{1}$ with (a) $g^{l}=0.125J_{1}$, (b) $g^{l}=0.3J_{1}$, (c) $g^{l}=0.586J_{1}$, (d) $g^{l}=0.8J_{1}$. The numbers highlighted in yellow denote the Chern numbers corresponding to the bands shown in (b) and (c).
Figure 5: (a) A combined phase diagram based on the Chern number sets of the corresponding bands is illustrated as a function of the non-local coupling $g^{nl}$ and the out-of-plane DMI $D$. The bulk band structures are shown along the high symmetry points with $B=0.4B_{s}$, $J_{2}=0.03J_{1}$, $D=0.045J_{1}$, $\hbar\omega=J_{1}S$ for (b) $g^{nl}=0.3J_{1}$, (c) $g^{nl}=0.438J_{1}$, and (d) $g^{nl}=0.8J_{1}$. The Chern numbers corresponding to the bands are denoted in yellow. These figures ascertain the occurrence of topological transitions upon solely tuning $g^{nl}$.
@@
Here, from Eq. (12), we observe that in the presence of local spin-phonon interaction ($H_{\text{sp}}^{l}$), the effective hopping strengths are reduced by a factor of $e^{-\lambda_{q}^{l}}$, where $q\in 1,2,3,4$ (see Eq. (33)), which helps to stabilize the low-energy magnon-polaron states. On the other hand, the onsite energy is modified by a negative energy shift $\Delta^{l}$ proportional to $(\alpha^{l})^{2}\omega$, which tends to destabilize the magnon-polaron states. Since our system already hosts a Goldstone mode near the $\mathbf{\Gamma}$ point, if the effect of $\Delta^{l}$ outweighs that of the $e^{-\lambda_{q}^{l}}$ term at large $\omega$, the system may slip into a regime with magnon-polaron instability, leading to imaginary eigenvalues within the BZ. Therefore, we keep $\hbar\omega=J_{1}S$, where the exponential factor dominates and hence avoids the magnon-polaron instability.
@@
In addition, to ensure that the system undergoes a topological phase transition, we calculate the Chern numbers corresponding to these bands and observe a transition from $C=3\rightarrow C=1$ for the lower band and from $C=-4\rightarrow C=-2$ for the middle band, in the presence of $g^{l}$ (at $g^{l}=0.2J_{1}$), as shown in Fig. 3(c). Hence, we observe that the presence of local spin-phonon coupling can significantly influence the topological state of the system. Furthermore, to examine the interplay between DMI and the spin-phonon coupling strength, we construct a topological phase diagram, shown in Fig. 3(d). Here, distinct topological regions are denoted by the Chern number set $C=(C_{1},C_{2},C_{3})$, where the indices $1$, $2$, and $3$ correspond to the lower, middle, and upper bands, respectively. Furthermore, in this phase diagram, we fix the other parameters at values given by $B=0.4B_{s}$, $J_{2}=0.03J_{1}$, and set the temperature to zero ($T=0$). Then, by assuring $D$ to be fixed and varying $g^{l}$ from $1$ to $J_{1}$, we observe multiple topological phase transitions along the $\text{u}_{1}$ and $\text{u}_{2}$ transition lines corresponding to a given classical ground state. Moreover, each of these transitions must be associated with a gap closing transition, as can be seen from spectral features.